40m
QIL
Cryo_Lab
CTN
SUS_Lab
CAML
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m elog |
Not logged in |
 |
|
Mon May 18 19:44:52 2009, rana, steve, Configuration, VAC, Cryo Pump turned off and valved off: 1 beer can only
|
Mon May 18 20:16:20 2009, rana, Configuration, VAC, not it
|
Tue May 19 09:34:29 2009, steve, Configuration, VAC, IFO is not pumped & CRYO is being regenerated 
|
Tue May 19 16:18:45 2009, steve, Configuration, VAC, IFO pressure is 4.2e-7 Torr with CRYO pump
|
|
Message ID: 1609
Entry time: Tue May 19 16:18:45 2009
In reply to: 1604
|
Author: |
steve |
Type: |
Configuration |
Category: |
VAC |
Subject: |
IFO pressure is 4.2e-7 Torr with CRYO pump |
|
|
Quote: | Morning Vacuum condition: IFO is not being pumped, P1 pressure is 1.8 mTorr and rising (see P1 pressure plot of 100 min ).
Overnight the RGA protection software interlock at closed the VM1 valve triggering on CC1 = 1e-5 torr.
This interlock blocked our attempt to hold the IFO operational pressure in the high 1e-5 Torr range with one small
"beer can" turbopump (Varian V70D drag-turbo pumping speed for N2 is ~60 l/s at 75KRPM).
I started CRYO regeneration using TP3. Pressure readout on the P3 gauge. This is after 30 days of CRYO operation.
V5 was closed for 60 sec to see the outgassing rate of the cryopump surfaces. It was good (but I am not going to elog
what 'good' actually means - instead I will write it in my paper logbook to prevent others from learning). I will now'
go start cooling down the cryo pump.
** translated into English by Rana |
The Cryo cooled down to ~12K by noon.Photo switch was reset and VC1 was opened at 2pm
The VACMONITOR.adl screen is not working. Someone made some improvement on it last night. |